OpenStudy (anonymous):

help pleasee!!!!!!!!!!! For which system of equations is (–2, 2) the solution? A. 3x – 5y = –16 –2x + 3y = 10 B. –4x + 2y = 6 2x – 3y = 10 C. 4x + 3y = 2 5x + 4y = –2 D. 6x + 5y = –2 –8x + 5y = –6

OpenStudy (anonymous):

A

OpenStudy (anonymous):

I agree ^

OpenStudy (anonymous):

For which system of equations is (–2, 3) the solution? A. 3x + 2y = 0 4x + 3y = –1 B. –3x + 4y = –6 2x + 5y = 11 C. –x + 7y = 16 5x – 3y = 16 D. –x + 7y = 23 5x + 3y = –1

OpenStudy (anonymous):

@612skies CAN YOU HELP ME OUT ON THIS ONE

OpenStudy (anonymous):

d